### v1

# Design Stability in Adaptive Experiments: Implications for Treatment Effect Estimation

Abstract
We study the problem of estimating the average treatment effect (ATE) under sequentially adaptive treatment assignment mechanisms. In contrast to classical completely randomized designs, we consider a setting in which the probability of assigning treatment to each experimental unit may depend on prior assignments and observed outcomes. Within the potential outcomes framework [1], we propose and analyze two natural estimators for the ATE: the inverse propensity weighted (IPW) estimator and an augmented IPW (AIPW) estimator. The cornerstone of our analysis is the concept of design stability, which requires that as the number of units grows, either the assignment probabilities converge, or sample averages of the inverse propensity scores and of the inverse complement propensity scores converge in probability to fixed, non-random limits. Our main results establish central limit theorems for both the IPW and AIPW estimators under design stability and provide explicit expressions for their asymptotic variances. We further propose estimators for these variances, enabling the construction of asymptotically valid confidence intervals. Finally, we illustrate our theoretical results in the context of Wei’s adaptive coin design [2] and Efron’s biased coin design [3], highlighting the applicability of the proposed methods to sequential experimentation with adaptive randomization.
Keywords: Average treatment effect; Sequential treatment assignment; Design stability; Adaptive designs; IPW estimator; AIPW estimator.
1 Introduction
Estimating the average treatment effect is a foundational problem in causal inference, especially when evaluating interventions in fields such as healthcare [4], education [5], public policy [6], development economics [7], and digital experimentation [8]. Traditional methods often assume simple randomized designs with independent and identically distributed (i.i.d.) units and fixed treatment assignment probabilities. However, many real-world experiments depart from this idealized setting: units often arrive sequentially, treatment assignments may adapt over time based on previous allocations, observed outcomes, or covariate information, and the study population is finite. Such sequential finite-population setups are common in adaptive clinical trials, online A/B testing, and policy evaluations. In these scenarios, adaptively assigning treatments can lead to complex dependencies among units, causing traditional ATE estimators to become biased or inefficient and undermining the applicability of conventional asymptotic results. Despite its practical importance, this setting remains methodologically less explored.
This paper develops a general framework for ATE estimation and inference under sequential designs in finite populations. We study the asymptotic behavior of the average treatment effect estimators under a class of sequential Bernoulli assignment mechanisms, in which the probability of assigning treatment to unit $i$ may depend on the observed history up to that point. Formally, $\mathbb{P}(K_{i}=1\mid\mathcal{F}_{i-1})=p_{i},$ $p_{i}\in\mathcal{F}_{i-1},$ where $K_{i}=1$ indicates that the $i$th unit is assigned to treatment, and $\mathcal{F}_{i-1}$ denotes the sigma-field generated by the treatment assignments and outcomes of the first (i1)(i-1) units. Two classic examples of such assignments that will be discussed in this paper are Wei’s adaptive coin design [2], and Efron’s biased coin design [3]. In the former, the assignment probabilities are expressed as a non-increasing function of the relative imbalance until the previous step, whereas in the latter design these probabilities take constant values that depend on the treatment-control imbalance up to the previous step (for example, fixed values like $\eta$, (1-$\eta$), $\tfrac{1}{2}\bigr)$.
Our contributions proceed in four parts. First, we begin by analyzing a standard estimator of the average treatment effect: the inverse propensity weighting (IPW) estimator, and propose an improvement by introducing a finite-population version of its augmented version (AIPW) that is commonly defined and used in model-based frameworks. Second, for both estimators, we establish central limit theorems under general sequential designs that satisfy a newly defined property called design stability. Third, under two different forms of design stability - strong and weak - we derive estimators of the asymptotic variances of the treatment effect estimators. These variance estimators, and the corresponding confidence intervals for the ATE, are conservative in that they are asymptotically positively biased, leading to asymptotic overcoverage of the confidence intervals. However, the biases vanish under certain forms of treatment effect homogeneity, yielding correct asymptotic coverage of the confidence intervals. Finally, we specialize these results to the two concrete experimental designs mentioned above, arguing that one of them (Wei’s adaptive design [2]) satisfies the strong design stability condition, whereas the other (Efron’s design [3]) satisfies the weak design stability condition.
The remainder of the paper is organized as follows. Section 2 reviews relevant prior work. Section 3 formally defines the problem, introduces the potential outcomes framework, describes the sequential assignment structure, and presents the estimators of interest. Section 4 presents the main theoretical results, including central limit theorems for adaptive designs and conservative asymptotic variance estimators for confidence interval construction. These results are then specialized in Section 5 to two widely used adaptive treatment assignment mechanisms. Sections 5.1 and 5.2 examine the stability of these designs and present simulation studies illustrating the finite-sample performance of the proposed estimators and supporting the theoretical findings. Section 6 concludes with a discussion and directions for future research. Proofs of all main and auxiliary results are provided in the supplementary material.
2 Related work
Causal inference in experimental settings is typically framed through two paradigms. The infinite-superpopulation perspective views study units as random draws from an underlying population, with randomness arising from the data-generating process. In contrast, the finite-population or design-based perspective treats the set of units as fixed, with uncertainty introduced solely through the experimental design. This latter view, combined with the potential outcomes formulation, traces back to [1], who conceptualized each unit’s treatment and control responses as fixed quantities and attributed randomness entirely to randomization. While classical asymptotic theory [[, e.g.,]]lehmann1999, vandervaart2000 is often aligned with the superpopulation framework, many applications, particularly randomized trials and survey sampling, are more naturally analyzed from the finite-population perspective [[, e.g.,]]fisher1935, cochran1977, rosenbaum2002, imbens2015.
The study of asymptotic normality in causal inference can be traced back to results on simple random sampling. Classical central limit theorems were established by [15][16][17], with convenient formulations presented in [18, 9]. These sampling-based central limit theorems can also be viewed as special cases of the more general results for rank statistics [19, 20, 21, 22]. Further foundational work includes the theory of U-statistics developed by [23] and the weak convergence results of [24], which laid the groundwork for modern asymptotic theory in survey sampling and experimental design. Because treatment and control groups in randomized experiments correspond to simple random samples from the finite set of experimental units, these sampling-based central limit theorems are directly applicable to the difference-in-means estimator of the average treatment effect. This connection underlies much of the early asymptotic justification in randomization-based causal inference [[, e.g.,]]Liu2014-kk, Ding2018-lr, Fogarty2018, li2017general.
In modern applications such as adaptive clinical trials [29], online A/B tests [30], and adaptive policy experiments [31], treatment assignments may evolve in response to interim data, violating the independence assumptions of static designs. Such sequential mechanisms introduce dependence across units, requiring martingale‐based CLTs [32, Chapter 3] in place of classical i.i.d. arguments.
In spite of the recent explosion of research on design-based finite population inference, to the best of our knowledge, rigorous theory for finite-population central limit theorems under general sequential general sequential Bernoulli assignments remains scarce. Recent explorations on inference of ATE from adaptive designs have been done in a setting where the potential outcomes for each experimental unit are assumed to follow an unknown probability distribution $\mathcal{P}$ and the ATE is defined as the difference of expectations of the potential outcomes with respect to $\mathcal{P}$. In this setting, [33] established asymptotic normality of the difference-in-means estimator under an adaptive Bernoulli allocation rule, and [34] extended these results to the augmented inverse probability weighting (AIPW) estimator [35]. However, this setting is different from design or randomization-based inference, where the uncertainty in the data (and consequently in the estimator) is induced solely by the act of randomization.
The present work addresses this gap, establishing central limit theorems for IPW and AIPW-inspired estimators in finite populations under broad sequential designs in a purely design-based inferential framework, where the potential outcomes are assumed fixed.
3 Problem Description
Consider a study with $N$ experimental units indexed by $i=1,\ldots,N$. We adopt the potential outcomes framework, introduced by [1] and later formalized by [36]. For each unit $i$, the outcome of interest $Y_{i}$ is characterized by two potential outcomes: $Y_{i}(0)$ under control and $Y_{i}(1)$ under treatment. The individual treatment effect is defined as $\tau_{i}=Y_{i}(1)-Y_{i}(0)$, and our target parameter is the average treatment effect (ATE), defined as
$\bar{\tau}=\frac{1}{N}\sum_{i=1}^{N}\tau_{i}\qquad\qquad(\textsc{ATE}).$ (1)
Assumptions of homogeneity of unit-level treatment effects $\tau_{1},\ldots,\tau_{N}$ play important roles in finite-population causal inference. For example, the assumption that the $\tau_{i}$’s are the same for $i=1,\ldots,N$, or equivalently,
$Y_{i}(1)=Y_{i}(0)+\tau\quad\text{for all }i,$ (2)
for some constant $\tau\in\mathbb{R}$ is called additivity of potential outcomes and is standard in literature [[, e.g.,]Chapter 6]imbens2015. Here we introduce the following definition that generalizes the concept of treatment effect homogeneity:
Definition 1 (Generalized treatment effect homogeneity).
Potential outcomes $(Y_{i}(0),Y_{i}(1))$, $i=1,\ldots,N$ are said to satisfy generalized treatment effect homogeneity if
$Y_{i}(1)-\overline{Y}_{N}(1)\propto Y_{i}(0)-\overline{Y}_{N}(0)\quad\text{for all }i,$ (3)
where $\overline{Y}_{N}(\ell)=\tfrac{1}{N}\sum_{i=1}^{N}Y_{i}(\ell)$ for $\ell\in\{0,1\}$.
It is easy to see that additivity (2) implies generalized treatment effect homogeneity (3). Another sufficient condition for (3) is additivity of potential outcomes on a log-scale, that is,
$Y_{i}(1)=cY_{i}(0)\quad\text{for all }i,$ (4)
for some constant $c\in\mathbb{R}$. We will see that conditions (2)-(4) play important roles in the inference problem to be discussed.
In the classical randomized treatment allocation design, a pre-defined constant number $N_{1}$ of the $N$ units are assigned to treatment, with the subset selected uniformly at random [36]. Formally, let $\mathbf{K}=(K_{1},K_{2},\ldots,K_{N})^{\text{T}}\in\{0,1\}^{N}$ denote the random assignment vector, where $K_{i}=1$ if $i$th unit is assigned to the treatment group and $K_{i}=0$ otherwise. A simple random sample of size $N_{1}$ is chosen from the finite population using the assignment vector $\mathbf{K}$, where $\mathbb{P}(\mathbf{K}=\mathbf{k})=\binom{N}{N_{1}}^{-1}$ for all $\mathbf{k}\in\{0,1\}^{N}$ satisfying $\mathbf{1}_{N}^{\text{T}}\mathbf{k}=N_{1}$. Given a treatment assignment vector $\mathbf{k}\in\{0,1\}^{N}$, the observed data $\{Y_{i}\}_{i=1}^{N}$ are the realized potential outcomes, where each unit’s outcome corresponds to its assigned treatment or control, defined by
$Y_{i}=K_{i}Y_{i}(1)+(1-K_{i})Y_{i}(0)\qquad\text{for}\qquad i=1,\ldots,N.$ (5)
A natural estimator of the ATE under the randomized treatment assignment described above is the difference in sample means between the treatment and control groups, i.e.
$\widehat{\widebar{\tau}}_{\text{avg}}$ $=\frac{1}{N_{1}}\sum_{i=1}^{N}K_{i}Y_{i}-\frac{1}{N_{0}}\sum_{i=1}^{N}(1-K_{i})Y_{i},\qquad(\textsc{difference-in-means estimator})$
where $N_{1}=\sum_{i=1}^{N}K_{i}\text{ and }N_{0}=N-N_{1}.$ This estimator is unbiased and satisfies a central limit theorem as $N$ grows to infinity [28]. We note that the difference-in-means estimator is a special case of the Horvitz–Thompson type estimator [37], also known as the inverse propensity weighted (IPW) estimator, defined as
$\widehat{\tau}_{\textsc{IPW}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left\{\frac{K_{i}Y_{i}}{p_{i}}-\frac{(1-K_{i})Y_{i}}{1-p_{i}}\right\},$ (6)
where the weights $p_{i}=N_{1}/N$ for $i=1,\ldots,N$.
In contrast to the classical randomized treatment assignment, we consider in this paper a sequential treatment assignment mechanism, in which the probability of assigning the $i$th unit to treatment is adaptive. In other words, for each unit $i\in{1,2,\ldots,N}$, the assignment indicator $K_{i}$, conditional on the past history, follows a Bernoulli distribution with success probability $p_{i}$, which we refer to as the inclusion probability. The inclusion probability $p_{i}$ is not fixed; rather, it is a measurable function of the prior assignment history and outcomes. Formally, we define a sequence of increasing sigma-fields $\{\mathcal{F}_{i-1}\}_{i\geq 1}$, where $\mathcal{F}_{i-1}=\sigma(K_{1},Y_{1},\ldots,K_{i-1},Y_{i-1})$ represents the cumulative information available after assigning treatment or control and observing the outcome of the (i1)(i-1)th unit. The assignment mechanism is such that
$p_{i}\in\mathcal{F}_{i-1}\quad\text{and}\quad\mathbb{P}(K_{i}=1\mid\mathcal{F}_{i-1})=p_{i}.\qquad(\textsc{Sequential treatment assignment})$ (7)
In words, the probability of assigning treatment to unit $i$ may depend on all previous assignments and observations up to stage (i1)(i-1) in an arbitrary way; we call this assignment a sequential treatment assignment. As an example, one may choose $p_{i}$ to promote relative balance between the numbers of treatment and control assignments. In this case, $p_{i}$ can be defined as the complement of the moving average of past assignments:
$p_{i}=1-\frac{\sum_{j=1}^{i-1}K_{j}}{i-1}\qquad\text{for \;\; $i\geq 2,$ \;\;\;\; and}\qquad p_{1}=\frac{1}{2}.$
In this assignment, if the previous units have mostly been assigned treatment, the chance of assigning treatment to the next unit will be lowered and vice versa [2].
In this paper, we aim to develop estimators, establish their asymptotics, and provide valid inference for the ATE under the sequential treatment assignment scheme (7). We first consider the Horvitz–Thompson type estimator defined in (6) as a natural unbiased estimator of the average treatment effect. In our setting, unlike a static completely randomized experiment, the $p_{i}$’s will not be equal and will depend on the past history.
It is well known that IPW estimators suffer from inflated variance when the probabilities approach extremes [38]. In a model-based setting, this limitation of IPW estimator is mitigated by the AIPW estimator [35] via model augmentation, offering double robustness. In our setting, where no probability model for the potential outcomes is assumed, we propose the following finite-population model-free version of the AIPW estimator:
$\widehat{\tau}_{\textsc{AIPW}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left[\left\{\frac{K_{i}(Y_{i}-\widehat{Y}_{i-1}(1))}{p_{i}}+\widehat{Y}_{i-1}(1)\right\}-\left\{\frac{(1-K_{i})(Y_{i}-\widehat{Y}_{i-1}(0))}{1-p_{i}}+\widehat{Y}_{i-1}(0)\right\}\right],$ (8)
where $\widehat{Y}_{1}(0)=\widehat{Y}_{1}(1)=0$, and for $i\geq 2$
$\widehat{Y}_{i-1}(0)=\frac{1}{i-1}\sum_{j=1}^{i-1}\frac{(1-K_{j})Y_{j}}{1-p_{j}},\qquad\text{and}\qquad\widehat{Y}_{i-1}(1)=\frac{1}{i-1}\sum_{j=1}^{i-1}\frac{K_{j}Y_{j}}{p_{j}}.$ (9)
Note that for large NN, the weighted average $\widehat{Y}_{N}(\ell)$ serves as an intuitive estimator of $\overline{Y}_{N}(\ell)$ for $\ell\in\{0,1\}$. In this sense, $\widehat{\tau}_{\textsc{AIPW}}$ is directly motivated from the classical AIPW estimator [35]. We formalize this intuition in a later theorem on the behavior of the AIPW estimator $\widehat{\tau}_{\textsc{AIPW}}$ (see Theorem 4).
4 Main Results
This section presents our main theoretical contributions. We establish the asymptotic normality of the estimators $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$ under the sequential experimental designs in (7), and derive conservative variance estimators that facilitate the construction of asymptotically valid confidence intervals for the average treatment effect $\bar{\tau}$. Our analysis proceeds in two steps: first, we prove central limit theorems for both estimators; second, we propose conservative estimators of their asymptotic variances. Together, these results enable the construction of asymptotically valid confidence intervals for $\bar{\tau}$.
The foundation of our analysis rests on a structural condition that we call design stability. We consider two notions of design stability (a) strong stability, and (b) weak stability. Strong design stability requires that the assignment probabilities themselves converge asymptotically, ensuring that the design does not drift in the limit. Weak design stability, however, relaxes this by requiring only that the sample averages of the inverse propensity scores and of the inverse complement propensity scores converge in probability to finite, non-random limits. At a high level, both forms of stability ensure that the cumulative effect of sequential randomization does not induce excessive variability in the long run.
Definition 2 (Strong design stability).
A sequential design with inclusion probabilities $\{p_{i}\}_{i\geq 1}$ is said to be strongly stable if there exists a non-random scalar $p^{\star}\in(0,1)$ such that
$p_{i}\xrightarrow{p}p^{\star}.$ (10)
Although the notion of strong design stability is intuitive, it is not satisfied by several popular designs. A concrete example is Efron’s biased coin design [3], which enforces balance between treatment and control assignments. Fortunately, Definition 2 can be relaxed so that, even if a design is not stable in the strong sense, central limit theorems for the IPW and AIPW estimators may still hold under weaker regularity conditions. This motivates the following weaker notion of stability.
Definition 3 (Weak design stability).
A sequential design with inclusion probabilities $\{p_{i}\}_{i\geq 1}$ is said to be weakly stable if there exists non-random scalars $p_{1}^{\star},p_{2}^{\star}\in(0,1)$ such that
$\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}\xrightarrow{p}\frac{1}{p_{1}^{\star}}\qquad\text{and}\qquad\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}\xrightarrow{p}\frac{1}{1-p_{2}^{\star}}.$ (11)
The tradeoff between these two stability notions becomes apparent when estimating the asymptotic variance of our estimators to construct confidence intervals for the ATE. While variance estimators can be constructed in a completely data-dependent manner under strong stability, weak stability requires additional restrictions (see Theorems 3 and 6 for more details).
4.1 The IPW estimator
We now turn to the asymptotic behavior of the IPW estimator $\widehat{\tau}_{\textsc{IPW}}$, as defined in (6). To derive our main result, we impose a positivity condition on the inclusion probabilities along with uniform boundedness and natural moment conditions on the potential outcomes.
Assumption 1.
The inclusion probabilities and potential outcomes satisfy the following regularity conditions:
  1. (a)
    There exists $\delta\in(0,1)$ such that $p_{i}\in[\delta,1-\delta]$ for all $i\geq 1$.
  2. (b)
    There exists a constant $M>0$ such that
    $|Y_{i}(\ell)|\leq M\quad\text{for all }i\geq 1\text{ and }\ell\in\{0,1\}.$
  3. (c)
    The following limits exist:
    $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}Y_{i}(0)^{2}=m^{2}_{0},\quad\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}Y_{i}(1)^{2}=m^{2}_{1},\quad\text{and}\quad\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}Y_{i}(0)Y_{i}(1)=m_{01},$
    where $m^{2}_{0},m^{2}_{1}>0$ and $m_{01}\in\mathbb{R}$.
The first condition in Assumption 1 ensures that the IPW estimator (6) is well defined, the second condition is a uniform bound on the potential outcomes, and the third assumption ensures that the limiting asymptotic variance of $\widehat{\tau}_{\textsc{IPW}}$ exists. With this set-up, we have the following guarantees on the asympotic behavior of $\widehat{\tau}_{\textsc{IPW}}$.
Theorem 1.
Suppose Assumption 1 holds, and the sequential design with inclusion probabilities $\{p_{i}\}_{i\geq 1}$ is either strongly or weakly stable in the sense of Definition 2 or Definition 3, respectively. Then the IPW estimator (6) satisfies
$\sqrt{N}\left(\widehat{\tau}_{\textsc{IPW}}-\bar{\tau}\right)\xrightarrow{d}\mathcal{N}\left(0,\mathrm{V}^{\mathrm{IPW}}\right),$ (12)
with asymptotic variance
$\mathrm{V}^{\mathrm{IPW}}=\begin{cases}\mathrm{V}^{\text{IPW}}_{\text{strong}}=m^{2}_{0}\dfrac{p^{\star}}{1-p^{\star}}+m^{2}_{1}\dfrac{1-p^{\star}}{p^{\star}}+2m_{01}&\text{(strong design stability)},\\[11.99998pt] \mathrm{V}^{\text{IPW}}_{\text{weak}}=m^{2}_{0}\dfrac{p_{2}^{\star}}{1-p_{2}^{\star}}+m^{2}_{1}\dfrac{1-p_{1}^{\star}}{p_{1}^{\star}}+2m_{01}&\text{(weak design stability)}.\end{cases}$ (13)
Remark 1.
The proof of Theorem 1 proceeds by rewriting the centered and scaled IPW estimator as a sum of martingale difference terms. This representation allows us to apply the martingale central limit theorem [32, Chapter 3]. We establish unbiasedness by verifying that each summand has zero conditional mean, and then compute the conditional variance, which converges under both stable and weakly stable designs to the stated asymptotic variance. Finally, we check the Lindeberg condition, ensuring that the contribution of large deviations vanishes. Together, these steps yield asymptotic normality of the IPW estimator with the asymptotic variance given in (13). Refer to Supplementary material 7.1 for detailed proof.
Having established the asymptotic normality of the IPW estimator, we next construct confidence intervals for the average treatment effect $\bar{\tau}$. This, in turn, requires estimation of the asymptotic variance $\mathrm{V}^{\mathrm{IPW}}$. First, we consider that the design is strongly stable in the sense of Definition 2, and assume that $p^{*}$ is known (which is the case in our illustrative example on strongly stable designs). To estimate $\mathrm{V}^{\text{IPW}}_{\text{strong}}$ in (13), we must estimate $m^{2}_{0}$, $m^{2}_{1}$, and $m_{01}$. While obtaining consistent estimators of $m^{2}_{0}$ and $m^{2}_{1}$ under strong stability is straightforward, the cross-moment term $m_{01}$ cannot be estimated from the observed outcomes without additional assumptions, as only one potential outcome is observed for each unit. To address this problem, we apply the Cauchy–Schwarz inequality to obtain $|m_{01}|\leq m_{0}m_{1}$, leading to the conservative variance estimator:
$\widehat{V^{\text{IPW}}_{\text{strong}}}=\left(\widehat{m}_{0}\sqrt{\frac{p^{\star}}{1-p^{\star}}}+\widehat{m}_{1}\sqrt{\frac{1-p^{\star}}{p^{\star}}}\right)^{2},$ (14)
where $\widehat{m}_{0}$ and $\widehat{m}_{1}$ are estimators of $m_{0}$ and $m_{1}$ that are consistent under strong stability. We propose the following intuitive estimators for $m^{2}_{0}$ and $m^{2}_{1}$
$\widehat{m}^{2}_{0}=\frac{1}{\max\{N_{0},1\}}\sum_{i=1}^{N}(1-K_{i})Y_{i}^{2},\qquad\text{and}\qquad\widehat{m}^{2}_{1}=\frac{1}{\max\{N_{1},1\}}\sum_{i=1}^{N}K_{i}Y_{i}^{2},\qquad$ (15)
where $N_{1}=\sum_{i=1}^{N}K_{i}$ and $N_{0}=N-N_{1}$.
The variance estimator in (14), which incorporates the estimators $\widehat{m}_{0}^{2}$ and $\widehat{m}_{1}^{2}$ defined in (15), is consistent when the potential outcomes are additive on a log-scale, that is, satisfy (4) and has an asymptotic positive bias otherwise. Hence, we obtain the following theorem.
Theorem 2.
For strongly stable designs (Definition 2), the estimators $\widehat{m}_{0}^{2}$ and $\widehat{m}_{1}^{2}$ defined in (15) are consistent for $m^{2}_{0}$ and $m^{2}_{1},$ respectively. Furthermore, the variance estimator $\widehat{V^{\text{IPW}}_{\text{strong}}}$ given by (14) provides a conservative estimate of $\mathrm{V}^{\text{IPW}}_{\text{strong}}$, and is consistent when the potential outcomes are additive on a log scale, that is, satisfy (4).
See Supplementary material 7.2 for a proof of the theorem.
Remark 2.
If for a strongly stable design the limiting value $p^{\star}$ is difficult to compute explicitly, the following consistent estimator
$\widehat{p}^{\star}=\frac{1}{N}\sum_{i=1}^{N}p_{i},\qquad$ (16)
may be substituted for $p^{*}$ into the variance estimator $\widehat{V^{\text{IPW}}_{\text{strong}}}$ (14).
We now turn to weakly stable designs (Definition 3) and assume that the limiting quantities $p_{1}^{\star}$ and $p_{2}^{\star}$ are known (which is the case in our illustrative example on weakly stable designs). Using arguments exactly analogous to the strongly stable case, we obtain the following conservative estimator of $\mathrm{V}^{\text{IPW}}_{\text{weak}}$ (13):
$\widehat{\mathrm{V}^{\text{IPW}}_{\text{weak}}}=\widetilde{m}_{0}^{2}{\frac{p_{2}^{\star}}{1-p_{2}^{\star}}}+\widetilde{m}_{1}^{2}{\frac{1-p_{1}^{\star}}{p_{1}^{\star}}}+2\widetilde{m}_{0}\widetilde{m}_{1},$ (17)
where $\widetilde{m_{0}}$ and $\widetilde{m_{1}}$ are estimators of $m_{0}$ and $m_{1}$ that are consistent under weak stability. However, unlike strongly stable designs, it is difficult to consistently estimate $m^{2}_{0}$ and $m^{2}_{1}$ without further restrictions. This illustrates the tradeoff between strong and weak design stability: although weak stability enlarges the class of admissible designs, it requires additional assumptions for consistent estimation of $m^{2}_{0}$ and $m^{2}_{1}$, and hence for conservative variance estimation. In particular, under the mild assumption (though not necessarily minimal) that, for some constant $\widetilde{p}\in(0,1)$,
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\widetilde{p}$, (18)
a consistent estimator of $m^{2}_{0}$ is
$\widetilde{m}_{0}^{2}=\frac{1}{N(1-\widetilde{p})}\sum_{i=1}^{N}(1-K_{i})Y_{i}^{2},$ (19)
with the analogous estimator for $m^{2}_{1}$ given by
$\widetilde{m}_{1}^{2}=\frac{1}{N\widetilde{p}}\sum_{i=1}^{N}K_{i}Y_{i}^{2}.$ (20)
As in the strongly stable case, the variance estimator in (17), which incorporates the estimators $\widetilde{m}_{0}^{2}$ and $\widetilde{m}_{1}^{2}$ defined in (19) and (20), respectively, is consistent when the potential outcomes are additive on the log-scale; that is, when they satisfy (4). Thus, we obtain the following theorem.
Theorem 3.
For weakly stable designs (Definition 3), under the sufficient condition (18), the estimators $\widetilde{m}_{0}^{2}$ and $\widetilde{m}_{1}^{2}$ from (19) and (20) are consistent for $m^{2}_{0}$ and $m^{2}_{1}$, respectively. Furthermore, the variance estimator $\widehat{V^{\text{IPW}}_{\text{weak}}}$ given by (17) provides a conservative estimate of $V^{\text{IPW}}_{\text{weak}}$, and is consistent when the potential outcomes are additive on a log scale, that is, satisfy (4).
Refer to Supplementary material 7.3 for a proof of the theorem.
Remark 3.
If, for a weakly stable design, the limiting values $p_{1}^{\star},p_{2}^{\star},$ and $\widetilde{p}$ are unknown or difficult to compute explicitly, we need to estimate them. Since $p_{i}\in\mathcal{F}_{i-1}$, the inclusion probability is deterministically known to the experimenter given the history. We therefore propose the following intuitive and consistent estimators for $p_{1}^{\star},p_{2}^{\star}$ and $\widetilde{p}$:
$\widehat{p}_{1}^{\star}=\frac{1}{\tfrac{1}{N}\sum_{i=1}^{N}\tfrac{1}{p_{i}}},\qquad\widehat{p}_{2}^{\star}=1-\frac{1}{\tfrac{1}{N}\sum_{i=1}^{N}\tfrac{1}{1-p_{i}}}\qquad\text{and}\qquad\overline{p}=\frac{1}{N}\sum_{i=1}^{N}p_{i}.$ (21)
Substituting the above estimators into (17), a conservative estimator of $\mathrm{V}^{\mathrm{IPW}}$ under weak stability when $p_{1}^{\star},p_{2}^{\star}$ and $\widetilde{p}$ are unknown is:
$\widehat{V^{\text{IPW}}_{\text{weak}}}=\widetilde{m}_{0}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}-1\right)+\widetilde{m}_{1}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}-1\right)+2\widetilde{m}_{0}\widetilde{m}_{1},$ (22)
where $\widetilde{m}_{0}^{2}=\displaystyle\frac{1}{\sum_{i=1}^{N}(1-p_{i})}\sum_{i=1}^{N}(1-K_{i})Y_{i}^{2}$ and $\widetilde{m}_{1}^{2}=\displaystyle\frac{1}{\sum_{i=1}^{N}p_{i}}\sum_{i=1}^{N}K_{i}Y_{i}^{2}.$
Since we have constructed conservative variance estimators of $\mathrm{V}^{\mathrm{IPW}}$, for any target level $\alpha\in(0,1)$, asymptotically conservative confidence interval for $\bar{\tau}$, that is, one with coverage at least $(1-\alpha)$ can be constructed,
$\lim_{N\rightarrow\infty}\mathbb{P}\left(\bar{\tau}\in\left[\widehat{\tau}_{\textsc{IPW}}-z_{1-\alpha/2}\cdot\sqrt{\frac{\widehat{V}}{N}},\;\;\widehat{\tau}_{\textsc{IPW}}+z_{1-\alpha/2}\cdot\sqrt{\frac{\widehat{V}}{N}}\right]\right)\geq 1-\alpha,$
where $\widehat{V}=\widehat{V^{\text{IPW}}_{\text{strong}}}$ or $\widehat{V^{\text{IPW}}_{\text{weak}}}$, and $z_{1-\alpha/2}$ denotes the $(1-\alpha/2)$th quantile of the standard normal distribution. Moreover, if the potential outcomes satisfy the log-additive treatment-effect model (4), the inequality holds with equality, yielding asymptotically exact coverage.
4.2 The AIPW-type estimator
We now analyze the asymptotic behavior of the AIPW estimator $\widehat{\tau}_{\textsc{AIPW}}$, as defined in (8).
Assumption 2.
The inclusion probabilities and potential outcomes satisfy the following regularity conditions:
  1. (a)
    There exists $\delta\in(0,1)$ such that $p_{i}\in[\delta,1-\delta]$ for all $i\geq 1$.
  2. (b)
    There exists a constant $M>0$ such that
    $|Y_{i}(\ell)|\leq M\quad\text{for all }i\geq 1\text{ and }\ell\in\{0,1\}.$
  3. (c)
    The following limits exist for $\ell\in\{0,1\}:$
    $\lim_{N\to\infty}\widebar{Y}_{N}(\ell)=\widebar{Y}_{\ell},\qquad\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}(\ell)-\widebar{Y}_{N}(\ell)\right)^{2}=\sigma^{2}_{\ell},$
    $\text{and}\qquad\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\left(Y_{i}(0)-\widebar{Y}_{N}(0)\right)\left(Y_{i}(1)-\widebar{Y}_{N}(1)\right)=\sigma_{01},$
    where $\sigma^{2}_{0},\sigma^{2}_{1}>0$ and $\widebar{Y}_{0},\widebar{Y}_{1},\sigma_{01}\in\mathbb{R}$.
The first two conditions in Assumption 2 are same as that of Assumption 1. The last condition above ensures that the limiting variance of the AIPW estimator exists. With this set-up, we state the asymptotic behavior of the AIPW estimator $\widehat{\tau}_{\textsc{AIPW}}$.
Theorem 4.
Suppose Assumption 2 holds, and the sequential design with inclusion probabilities $\{p_{i}\}_{i\geq 1}$ is either strongly or weakly stable in the sense of Definition 2 or Definition 3, respectively. Then the AIPW estimator (8) satisfies
$\sqrt{N}\left(\widehat{\tau}_{\textsc{AIPW}}-\bar{\tau}\right)\xrightarrow{d}\mathcal{N}\left(0,\mathrm{V}^{\mathrm{AIPW}}\right),$ (24)
with asymptotic variance
$\mathrm{V}^{\mathrm{AIPW}}=\begin{cases}\mathrm{V}^{\text{AIPW}}_{\text{strong}}=\sigma^{2}_{0}\dfrac{p^{\star}}{1-p^{\star}}+\sigma^{2}_{1}\dfrac{1-p^{\star}}{p^{\star}}+2\sigma_{01}&\text{under strong design stability},\\[11.99998pt] \mathrm{V}^{\text{AIPW}}_{\text{weak}}=\sigma^{2}_{0}\dfrac{p_{2}^{\star}}{1-p_{2}^{\star}}+\sigma^{2}_{1}\dfrac{1-p_{1}^{\star}}{p_{1}^{\star}}+2\sigma_{01}&\text{under weak design stability}.\end{cases}$ (25)
Remark 4.
The proof of Theorem 4 requires a different strategy from that of Theorem 1, since the AIPW estimator is not directly amenable to martingale central limit theorem. To handle this, we introduce a proxy estimator that is analytically more tractable and can be expressed as sum of a martingale difference sequence, allowing the martingale central limit theorem to establish its asymptotic normality. The key step is then to show that the difference between the proxy and the actual AIPW estimator is asymptotically negligible, using variance bounds and Hájek’s lemma (see Supplementary material 6). This ensures that the asymptotic distribution of the AIPW estimator coincides with that of the proxy, yielding the stated central limit theorem with variance given in Theorem 4. See Supplementary material 7.4 for detailed proof of the above theorem.
Remark 5.
Note that $\mathrm{V}^{\mathrm{AIPW}}\leq\mathrm{V}^{\mathrm{IPW}}$; that is, $\widehat{\tau}_{\textsc{AIPW}}$ is more efficient than $\widehat{\tau}_{\textsc{IPW}}$. This fact clearly establishes the superiority of the AIPW estimator over the IPW estimator in finite population design-based inference under the adaptive assignment mechanism defined in (7).
We now turn to the problem of estimating the asymptotic variance $\mathrm{V}^{\mathrm{AIPW}}$. Under strong design stability (Definition 2) with known $p^{\star}$, estimation of $\mathrm{V}^{\text{AIPW}}_{\text{strong}}$ in (25) requires estimation of $\sigma^{2}_{0}$, $\sigma^{2}_{1}$ and $\sigma_{01}$. As earlier, the covariance term $\sigma_{01}$ depends on both potential outcomes for the same unit and therefore cannot be estimated without additional restrictions. Analogous to the estimation of $\mathrm{V}^{\mathrm{IPW}}$ in Section 4.1, we invoke Cauchy-Schwarz inequality to get $|\sigma_{01}|\leq\sigma_{0}\sigma_{1}$, yielding the following conservative estimator of $\mathrm{V}^{\text{AIPW}}_{\text{strong}}$ as follows:
$\widehat{V^{\text{AIPW}}_{\text{strong}}}=\left(\widehat{\sigma}_{0}\sqrt{\frac{p^{\star}}{1-p^{\star}}}+\widehat{\sigma}_{1}\sqrt{\frac{1-p^{\star}}{p^{\star}}}\right)^{2},$ (26)
where $\widehat{\sigma}^{2}_{0}$ and $\widehat{\sigma}^{2}_{1}$ are estimators of $\sigma_{0}$ and $\sigma_{1}$ that are consistent under strong design stability. We propose the following estimators:
$\widehat{\sigma}^{2}_{0}$ $=\frac{1}{\max\{N_{0},1\}}\sum_{i=1}^{N}(1-K_{i})\big(Y_{i}-\widehat{Y}_{i-1}(0)\big)^{2},$ (27a)
$\widehat{\sigma}^{2}_{1}$ $=\frac{1}{\max\{N_{1},1\}}\sum_{i=1}^{N}K_{i}\big(Y_{i}-\widehat{Y}_{i-1}(1)\big)^{2},$ (27b)
where $N_{1}=\sum_{i=1}^{N}K_{i}$ and $N_{0}=N-N_{1}$. We set $\widehat{Y}_{1}(0)=\widehat{Y}_{1}(1)=0$, and for $i\geq 2$ define
$\widehat{Y}_{i-1}(0)=\frac{1}{i-1}\sum_{j<i}\frac{(1-K_{j})Y_{j}}{1-p_{j}},\qquad\widehat{Y}_{i-1}(1)=\frac{1}{i-1}\sum_{j<i}\frac{K_{j}Y_{j}}{p_{j}}.$
The variance estimator in (26), which incorporates the estimators in (27), is consistent when the potential outcomes satisfy the generalized treatment effect homogeneity condition in Definition 1. The preceding discussion leads to the following theorem.
Theorem 5.
For strongly stable designs (Definition 2), the estimators $\widehat{\sigma}^{2}_{0}$ and $\widehat{\sigma}_{1}^{2}$ defined in (27) are consistent for $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$, respectively. Furthermore, the variance estimator $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ given by (26) provides a conservative estimate of $\mathrm{V}^{\text{AIPW}}_{\text{strong}}$, and is consistent when the potential outcomes satisfy generalized treatment effect homogeneity (3).
A proof of this theorem is given in Supplementary material 7.5.
As noted in Section 4.1, under weak design stability (Definition 3) with known $p_{1}^{\star}$ and $p_{2}^{\star}$, variance estimation is not straightforward, as additional conditions are required for the consistent estimation of $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$. As before, under the additional assumption (18) and with known $\widetilde{p}$, the variance components $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$ can be consistently estimated by:
$\widetilde{\sigma}^{2}_{0}$ $=\frac{1}{N(1-\widetilde{p})}\sum_{i=1}^{N}(1-K_{i})\big(Y_{i}-\widehat{Y}_{i-1}(0)\big)^{2},$ (28)
$\widetilde{\sigma}^{2}_{1}$ $=\frac{1}{N\widetilde{p}}\sum_{i=1}^{N}K_{i}\big(Y_{i}-\widehat{Y}_{i-1}(1)\big)^{2}.$ (29)
As discussed previously, the cross-moment term $\sigma_{01}$ cannot be estimated without additional assumptions, since it depends on both potential outcomes for all units. We therefore bound it from above using the Cauchy–Schwarz inequality. Consequently, asymptotic variance $\mathrm{V}^{\text{AIPW}}_{\text{weak}}$ can be conservatively estimated by:
$\widehat{V^{\text{AIPW}}_{\text{weak}}}=\widetilde{\sigma}_{0}^{2}{\frac{p_{2}^{\star}}{1-p_{2}^{\star}}}+\widetilde{\sigma}_{1}^{2}{\frac{1-p_{1}^{\star}}{p_{1}^{\star}}}+2\widetilde{\sigma}_{0}\widetilde{\sigma}_{1}.$ (30)
As in the strong stability case, this estimator is consistent when the potential outcomes satisfy generalized treatment effect additivity according to Definition 1. The above discussion is summarized in the following theorem.
Theorem 6.
For weakly stable designs (Definition 3), under the sufficient condition (18), the estimators $\widetilde{\sigma}_{0}^{2}$ and $\widetilde{\sigma}_{1}^{2}$ from (28) and (29) are consistent for $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$, respectively. Furthermore, the variance estimator $\widehat{V^{\text{AIPW}}_{\text{weak}}}$ given by (30) provides a conservative estimate of $\mathrm{V}^{\text{AIPW}}_{\text{weak}}$, and is consistent when the potential outcomes satisfy generalized treatment effect additivity (3).
See Supplementary material 7.6 for the proof.
Remark 6.
If for a strongly stable design the limiting value $p^{\star}$ is unknown or difficult to compute explicitly, substitution of the consistent estimator $\widehat{p}^{\star}$ defined in (16) in place of $p^{\star}$ into (26) will lead to an estimator of $\mathrm{V}^{\text{AIPW}}_{\text{strong}}$ with similar properties as in Theorem 5. If for a weakly stable design the limiting values $p_{1}^{\star},p_{2}^{\star}$ and $\widetilde{p}$ are unknown or are difficult to compute explicitly, we can estimate them using (21). Substituting these estimators into (30), a conservative estimator of $\mathrm{V}^{\text{IPW}}_{\text{weak}}$ under weak stability is:
$\widehat{V^{\text{AIPW}}_{\text{weak}}}=\widetilde{\sigma}_{0}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{1-p_{i}}-1\right)+\widetilde{\sigma}_{0}^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{1}{p_{i}}-1\right)+2\widetilde{\sigma}_{0}\widetilde{\sigma}_{1},$ (31)
where $\widetilde{\sigma}_{0}^{2}=\displaystyle\frac{1}{\sum_{i=1}^{N}(1-p_{i})}\sum_{i=1}^{N}(1-K_{i})\big(Y_{i}-\widehat{Y}_{i-1}(0)\big)^{2}$ and $\widetilde{\sigma}_{1}^{2}=\displaystyle\frac{1}{\sum_{i=1}^{N}p_{i}}\sum_{i=1}^{N}K_{i}\big(Y_{i}-\widehat{Y}_{i-1}(1)\big)^{2}.$
As in Section 4.1, for any target level $\alpha\in(0,1)$, asymptotically conservative confidence intervals for $\bar{\tau}$ may be constructed around $\widehat{\tau}_{\textsc{AIPW}}$ using the variance estimators (26) and (30) for strongly stable and weakly stable designs, respectively. If the limiting values of the probabilities are unknown, their counterparts suggested in Remark 6 may be used. All of these intervals are asymptotically conservative, but attain exact asymptotic coverage when the potential outcomes satisfy generalized treatment effect additivity (3).
5 Some illustrative applications
In this section, we illustrate Theorems 1-6 through two adaptive designs: a strongly stable design, Wei’s adaptive coin design [2], and a weakly stable design, Efron’s biased coin design [3]. We further complement the theoretical results with numerical simulations that demonstrate the validity of our approach.
5.1 Wei’s Adaptive Coin Design
We begin with Wei’s adaptive coin design [2], which reduces relative imbalance between treatment and control allocations. Formally, let $m_{k}$ and $n_{k}$ denote, respectively, the numbers of treatment and control units among the first $k$ subjects. Define the treatment-control imbalance as $D_{k}=m_{k}-n_{k},$ and the corresponding normalized imbalance $R_{k}=\frac{D_{k}}{k},$ which measures the average difference between the treatment and control groups up to stage $k$.
Under Wei’s adaptive coin design, the $i$th subject is assigned to treatment with probability
$p_{i}=f\left(R_{i-1}\right),$ (32)
where $f:[-1,1]\to[0,1]$ is a non-increasing function satisfying (i) $f(0)=\frac{1}{2}$ and (ii) $f$ is continuous at zero. For the estimators $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$ to be well-defined under this design, it is necessary that the inclusion probabilities be bounded away from zero and one. If $f$ does not guarantee this property, we may enforce it by replacing $p_{i}$ in (32) with the clipped version
$p_{i}=\min\big\{\max\{f(R_{i-1}),\delta\},\,1-\delta\big\},$ (33)
for some fixed $\delta\in(0,\frac{1}{2}]$. This modification ensures $p_{i}\in[\delta,\,1-\delta]$ for all $i\geq 1$.
Intuitively, when the trial is in its early stages, the number of units in each group can differ substantially in relative terms; the design then shifts $p_{i}$ away from $\tfrac{1}{2}$ to favor the smaller group and reduce imbalance. As the sample size grows, any absolute difference in group sizes becomes small relative to the total number of units, causing $R_{i-1}$ to shrink and $p_{i}$ to converge to $\tfrac{1}{2}$. The following Lemma, which is a direct consequence of [2, Theorem 1], establishes strong stability of the truncated version of Wei’s design (33), making Theorems 1245 directly applicable.
Lemma 1.
Wei’s adaptive coin design ($3$) is strongly stable in sense of Definition 2, with $p^{\star}=\frac{1}{2}$.
See Supplementary material 8.1 for the proof of Lemma 1.
Substituting $p^{\star}=\tfrac{1}{2}$ into the expressions for $\mathrm{V}^{\mathrm{IPW}}$ and $\mathrm{V}^{\mathrm{AIPW}}$ in Theorems 1 and 4 gives the limiting variances of the IPW and AIPW estimators, $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$, respectively, under Wei’s design:
$V^{\text{IPW}}_{\text{Wei}}$ $=m^{2}_{0}+m^{2}_{1}+2m_{01},$ (34a)
$V^{\text{AIPW}}_{\text{Wei}}$ $=\sigma^{2}_{0}+\sigma^{2}_{1}+2\sigma_{01}.$ (34b)
Since $p^{\star}$ is known and fixed at $\tfrac{1}{2}$, it can be directly plugged into the variance estimators $\widehat{V^{\text{IPW}}_{\text{strong}}}$ in (14) and $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ in (26), yielding the following conservative estimators for the IPW and AIPW variances:
$\widehat{V^{\text{IPW}}_{\text{Wei}}}=(\widehat{m}_{0}+\widehat{m}_{1})^{2},\qquad\widehat{V^{\text{AIPW}}_{\text{Wei}}}=(\widehat{\sigma}_{0}+\widehat{\sigma}_{1})^{2},$ (35)
where $\widehat{m}^{2}_{\ell}$ and $\widehat{\sigma}^{2}_{\ell}$, $\ell\in\{0,1\}$, are as defined in (15) and (27), respectively. We can now use these variance estimators in place of $\widehat{V^{\text{IPW}}_{\text{strong}}}$ and $\widehat{V^{\text{AIPW}}_{\text{strong}}}$ to construct conservative confidence intervals for $\bar{\tau}$. Recall that the interval based on $\widehat{V^{\text{IPW}}_{\text{Wei}}}$ attains exact asymptotic coverage when the potential outcomes satisfy additivity on the log scale (4), whereas the interval based on $\widehat{V^{\text{AIPW}}_{\text{Wei}}}$ attains exact asymptotic coverage under generalized treatment effect additivity, as defined in (3).
(a) IPW estimator under Non-Additive case
(b) AIPW estimator under Non-Additive case
(c) IPW estimator under Additive case
(d) AIPW estimator under Additive case
(e) IPW estimator under Log-Additive case
(f) AIPW estimator under Log-Additive case
Figure 1: Comparison of the theoretical and empirical coverages for Wei’s design.
(a) Non-Additive case
(b) Additive case
(c) Log-Additive case
Figure 2: Comparison of the average lengths of confidence intervals for Wei’s design.
Next, we evaluate the performances of the IPW and AIPW estimators under Wei’s adaptive coin design through simulation studies. We consider three data-generating mechanisms: (a) a general, non-additive outcome model; (b) the additive model in equation (2); and (c) the log-additive model in equation (4).
In the non-additive setting, the potential outcomes $(Y_{i}(0),Y_{i}(1))$ are drawn from a bivariate normal distribution with mean vector $(0,1)^{\text{T}}$ and variance–covariance matrix
$\begin{bmatrix}1&0.3\\ 0.3&1\end{bmatrix}$,
with support restricted to [3,3][-3,3] to ensure bounded outcomes.
In the additive setting, the control potential outcomes are drawn from a normal distribution with mean 0 and variance 1, truncated to [3,3][-3,3], and the treatment outcomes are defined by $Y_{i}(1)=Y_{i}(0)+\tau$ with $\tau=10$.
In the log-additive setting, the control potential outcomes are drawn from a normal distribution with mean 10 and variance 1, truncated to $[7,13]$, and the treatment outcomes are defined by $Y_{i}(1)=c\,Y_{i}(0)$ with $c=2$.
Treatment assignments $\mathbf{K}$ are generated according to Wei’s sequential randomization scheme, with assignment probabilities $p_{i}=f(R_{i-1})=\tfrac{1-R_{i-1}}{2}$, where $R_{i-1}$ denotes the normalized treatment–control imbalance prior to assigning the $i$th unit, and the truncation parameter is set to $\delta=0.01$. Each simulation involves N=5, [^0] N=5{,}000 units and is repeated 20,00020{,}000 times. For each replication, confidence intervals are constructed using the proposed methodology, and empirical coverage is evaluated across 20 nominal levels ranging from 0.75 to 0.99.
Figure 1 reports the empirical coverage of confidence intervals based on the IPW and AIPW estimators. In both cases, the intervals exhibit reliable coverage of the true $\bar{\tau}$. For the non-additive setup, both intervals remain conservative, whereas under additivity and log-additivity, the empirical coverage approaches the nominal levels. In particular, the AIPW estimator performs better in the non-additive setting, yielding coverage closer to the nominal levels than the IPW estimator. The IPW estimator attains nearly exact coverage under the log-additive setup, while the AIPW estimator achieves nearly exact coverage under additivity and remains close to nominal levels under log-additivity. These results align with the theoretical guarantees of variance estimator consistency established in Theorems 2 and 5, and overall demonstrate the superior stability of the AIPW estimator.
Figure 2 displays the average confidence interval lengths for the true parameter $\bar{\tau}$ under the IPW and AIPW estimators. Each interval length is computed as $2\,z_{1-\alpha/2}\,\sqrt{\widehat{V}/N},$ where $z_{1-\alpha/2}$ is the standard normal quantile corresponding to the nominal level $\alpha$, and $\widehat{V}$ denotes the estimated variance. For the IPW and AIPW estimators, this corresponds to $\widehat{V^{\text{IPW}}_{\text{Wei}}}$ and $\widehat{V^{\text{AIPW}}_{\text{Wei}}}$, respectively, as defined in equation (35). Across all confidence levels and data-generating mechanisms, the AIPW estimator produces substantially shorter intervals than the IPW estimator. Together with the coverage results established in Figure 1, these results highlight the overall greater efficiency and stability of the AIPW estimator. Specifically, while the IPW estimator achieves valid coverage under log-additive setup, the AIPW estimator performs remarkably better in both additive and non-additive setups, providing coverage levels closer to the nominal values along with consistently shorter confidence intervals. In general, these results emphasize that although both estimators attain reliable coverage, the AIPW estimator achieves this with noticeably tighter intervals, making it generally more efficient and preferable in practical applications.
5.2 Efron’s Biased Coin Design
Moving beyond Wei’s adaptive coin design, we consider the biased coin design introduced by [3] which enforces another form of approximate balance between the number of allocations in the treatment and control groups. As in the previous section, let $D_{k}=m_{k}-n_{k}$ denote the imbalance between the treatment and control groups after the assignment of the $k$th unit, where $m_{k}$ and $n_{k}$ denote the numbers of treatment and control assignments, respectively. Under Efron’s biased coin design, the $i$th unit is assigned to treatment with probability:
$p_{i}=\begin{cases}\eta&\text{if }D_{i-1}<0\\ \frac{1}{2}&\text{if }D_{i-1}=0\\ 1-\eta&\text{if }D_{i-1}>0\end{cases},$ (36)
where $\eta\in[\frac{1}{2},1)$ controls the strength of the bias toward balance. In words, a larger value of $\eta$ forces faster correction of imbalance.
It is worth noting that $p_{i}$ takes the values $\eta$ and $1-\eta$ infinitely often, and thus Efron’s biased coin design is not strongly stable in the sense of Definition 2. However, the following lemma establishes weak stability of the design, making Theorems 1, 3, 4, and 6 applicable.
Lemma 2.
Efron’s biased coin design (36) is weakly stable in sense of Definition 3, with $p_{1}^{\star}=\frac{4\eta^{2}(1-\eta)}{1-4\eta+12\eta^{2}-8\eta^{3}}$ and $p_{2}^{\star}=\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{1-4\eta+12\eta^{2}-8\eta^{3}}$. Moreover,
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\frac{1}{2}.$
Remark 7.
The proof of Lemma 2 proceeds by studying the treatment–control imbalance sequence $\{D_{k}\}_{k\geq 1}$. We first establish that $\{D_{k}\}_{k\geq 1}$ forms an irreducible and positively recurrent Markov chain by applying Foster’s Theorem [39]. The resulting positive recurrence and irreducibility ensure the existence of a unique stationary distribution, which, together with the mean ergodic theorem, facilitates characterization of the limiting behavior of long-run averages of functions of the assignment probabilities. Details of the proof are provided in Supplementary material 8.2.
Substituting the values of $p_{1}^{\star}$ and $p_{2}^{\star}$ from Lemma 2 into Theorems 1 and 4 yields the limiting variances of the IPW and AIPW estimators, $\widehat{\tau}_{\textsc{IPW}}$ and $\widehat{\tau}_{\textsc{AIPW}}$, under Efron’s design:
$V^{\text{IPW}}_{\text{Efron}}$ $=\left(m^{2}_{0}+m^{2}_{1}\right)\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{4\eta^{2}(1-\eta)}+2m_{01},$ (37a)
$V^{\text{AIPW}}_{\text{Efron}}$ $=\left(\sigma^{2}_{0}+\sigma^{2}_{1}\right)\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{4\eta^{2}(1-\eta)}+2\sigma_{01}.$ (37b)
Furthermore, as Lemma 2 shows, under this design $N^{-1}\sum_{i=1}^{N}p_{i}\xrightarrow{p}\tfrac{1}{2}$, satisfying the sufficient condition in (18). This result allows for consistent estimation of $m^{2}_{0}$ and $m^{2}_{1}$ from (19) and (20), and of $\sigma^{2}_{0}$ and $\sigma^{2}_{1}$ from (28) and (29). Substituting these estimates into (17) and (30) yields conservative estimators of $V^{\text{IPW}}_{\text{Efron}}$ (37a) and $V^{\text{AIPW}}_{\text{Efron}}$ (37b):
$\widehat{V^{\text{IPW}}_{\text{Efron}}}$ $=\left(\widehat{m}^{2}_{0}+\widehat{m}^{2}_{1}\right)\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{4\eta^{2}(1-\eta)}+2\widehat{m}_{0}\widehat{m}_{1},$ (38a)
$\widehat{V^{\text{AIPW}}_{\text{Efron}}}$ $=\left(\widehat{\sigma}^{2}_{0}+\widehat{\sigma}^{2}_{1}\right)\frac{1-4\eta+8\eta^{2}-4\eta^{3}}{4\eta^{2}(1-\eta)}+2\widehat{\sigma}_{0}\widehat{\sigma}_{1},$ (38b)
(a) IPW estimator under Non-Additive case
(b) AIPW estimator under Non-Additive case
(c) IPW estimator under Additive case
(d) AIPW estimator under Additive case
(e) IPW estimator under Log-Additive case
(f) AIPW estimator under Log-Additive case
Figure 3: Comparison of the theoretical and empirical coverages for Efron’s design.
(a) Non-Additive case
(b) Additive case
(c) Log-Additive case
Figure 4: Comparison of the average lengths of confidence intervals for Efron’s design.
As in Section 5.1, the variance estimators $\widehat{V^{\text{IPW}}_{\text{Efron}}}$ and $\widehat{V^{\text{AIPW}}_{\text{Efron}}}$ can be used to construct conservative confidence intervals for $\bar{\tau}$. The interval based on $\widehat{V^{\text{IPW}}_{\text{Efron}}}$ achieves exact asymptotic coverage when the potential outcomes satisfy additivity on the log scale (4), while the interval based on $\widehat{V^{\text{AIPW}}_{\text{Efron}}}$ attains exact coverage under generalized treatment effect additivity, as defined in (3).
We now assess the coverage of the confidence intervals constructed using the IPW and AIPW estimators under Efron’s design. The simulations use the same data-generating procedures and parameter settings as in Section 5.1, except that treatment assignments now follow Efron’s biased coin design (36) rather than Wei’s design. The biased-coin parameter is fixed at $\eta=0.7$ in all simulations. Figure 3 shows the empirical coverage of confidence intervals for the IPW and AIPW estimators, while Figure 4 reports the corresponding average interval lengths. The lengths are computed as for Wei’s design, using $\widehat{V^{\text{IPW}}_{\text{Efron}}}$ for IPW and $\widehat{V^{\text{AIPW}}_{\text{Efron}}}$ for AIPW, as specified in equations (38a) and (38b), respectively.
The results for Efron’s biased coin design are consistent with those observed under Wei’s adaptive design, as shown in Figures 1 and 2. All simulations were conducted for a population size of N=5, [^0] N=5{,}000, which is sufficiently large for the asymptotic approximations to apply in sequential experimental settings. Accordingly, the empirical findings align closely with the theoretical results established in Section 4. As established theoretically, the IPW estimator empirically attains nearly exact asymptotic coverage under the log-additive setup. The AIPW estimator attains nearly exact asymptotic coverage in the additive case and consistently remains closer to nominal levels than the IPW estimator in the non-additive setting, in agreement with theoretical expectations. Under the log-additive setup, the AIPW estimator also provides near-exact asymptotic coverage, matching the nominal reference line and confirming the consistency of its variance estimator. Across all scenarios, the AIPW estimator yields shorter confidence intervals than IPW, highlighting its overall efficiency. These findings corroborate the theoretical results stated in Theorems 2 and 5, and indicate that the AIPW estimator is more stable and efficient under both strong and weak design stabilities.
6 Discussion
We have developed a general theoretical framework for conducting inference on average treatment effects in settings where treatment assignment is sequentially adaptive within a finite population. This framework unifies and extends existing results by accommodating a broad class of adaptive randomization schemes, where assignment probabilities may evolve over time based on past outcomes. Within this setup, we establish central limit theorems (CLTs) for both inverse probability weighted (IPW) and augmented IPW (AIPW) estimators under strong and weak design stability conditions. Although the limiting distributions feature explicit expressions for the asymptotic variances, the fundamental problem of causal inference - not being able to observe the two potential outcomes for each unit - leads to challenges in their estimation. We propose conservative variance estimators that are consistent under different forms of treatment effect homogeneity.
To demonstrate the applicability of our framework, we analyze Wei’s adaptive coin design and Efron’s biased coin design, two classical examples in sequential experimentation. These applications reveal how the general theory accommodates designs that deviate from strong stability (e.g., Efron’s design), thereby illustrating its flexibility and robustness.
From a practical standpoint, our findings provide reassurance that adaptive treatment assignment mechanisms—increasingly popular in modern experimental and clinical trial settings—can be used within a finite population framework without imposing any model on the potential outcome. The research opens up several new research possibilities. Extending the framework to covariate-adaptive designs where assignments depend explicitly on pre-measured covariates [[, e.g.,]]FSM1979, morgan2012, would broaden the applicability of the theory. Adaptive treatment assignment mechanisms also provide a natural solution to finding optimal designs in a finite population setting, e.g., [42] and the results presented in this paper can provide an inferential framework for such adaptive designs.